If the velocity of hydrogen molecule is 5×100000cm/s then its de broglie wave length is

Grade:11

1 Answers

Harshit Panwar
6 years ago
de Broglie wave length :-
Wavelength = h/mv
=6.626 ×10-34 /0.001 kg × 5000 m/s
=6.626 × 10-34/5
=1.325 × 10-37 m
Hope it helps you.
Approve if it helps you.